Angus invested $18000, part at 3% and part at 5%. If the total interest at the end of the year is $660, how much did he invest at each rate?
Answer by checkley71(8403)

.05X+.03(18000-X)=660
.05X+540-.03X=660
.05X-.03X=660-540
.02X=120
X=120/.02
X=6000 AMOUNT INVESTED @ 5%.
18000-6000=12000 INVESTED @ 3%.
PROOF
.05*6000+.03*12000=660
300+360=660
660=660